# appendixReferences

### v3

@@ A.1 Auxilliary Results
Lemma 1 is used to prove Theorem 4. It applies McDiarmid’s inequality to show concentration of the terms appearing in $\hat{\Phi}_{k}^{\textup{Haj}}$ for $k=1,2$.
Lemma 1.
@@
Let Assumption 1 hold. For $i\in[N]$ and $\bar{Z}_{i}$ given by (39), let $u_{i}$ denote the vector
and let $\bar{u}=\frac{1}{N}\sum_{i=1}^{N}u_{i}$. Then $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}(N^{-1/2})$.
In our proof of Theorem 5, we will use $C_{ij}(T_{i},T_{j})$ to abbreviate the covariance estimates appearing in $V_{k}$. It is given by
Lemmas 2 and 3 are used to prove (47) of Theorem 5. Lemma 2 bounds the values of $\{C_{ij}\}$.
Lemma 2.
Let Assumption 1 hold and let $C_{ij}$ be given by (54). Let $\mathcal{I}_{0}$ and $\mathcal{I}_{1}$ denote the sets
$\mathcal{I}_{0}$ =$=\{i,j:\phi_{i}^{*}\phi_{j}^{*}=1,|\Gamma_{i}\cap\Gamma_{j}|=0\}$
$\mathcal{I}_{1}$ =$=\{i,j:\phi_{i}^{*}\phi_{j}^{*}=1,|\Gamma_{i}\cap\Gamma_{j}|\neq 0\}$
It holds w.p. 1 that
$\max_{i,j\in\mathcal{I}_{0}}C_{ij}(T_{i},T_{j})\leq O\left(\frac{\log N}{N}\right)$ (55)
$\max_{i,j\in\mathcal{I}_{1}}C_{ij}(T_{i},T_{j})\leq 1+O\left(\frac{\log N}{N}\right)$ (56)
To bound the size of the IPW weights appearing in the variance estimates, Lemma 3 states that certain events, which depend on only a finite number of treatments, cannot have probability that is infinitesimally small yet positive.
Lemma 3.
Let Assumption 1 hold, and let $\mathcal{T}_{i}$ denote the support of $T_{i}$. Then there exists a fixed constant $\rho_{\min}>0$ such that
$P(T_{i}=t_{i},T_{j}=t_{j})$$\in\{0\}\cap[\rho_{\min},1]\qquad\forall\ i,j\in[N]^{2}$ (57)
$P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})$$\in\{0\}\cap[\rho_{\min},1]\qquad\forall\ i,j:\phi_{i}^{*}\phi_{j}^{*}=1$ (58)
Lemmas 4 and 5 are used to prove (48) of Theorem 5.
Lemma 4.
Let $C_{ij}$ be given by (54). If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then $C_{ij}(T_{i},T_{j})\leq 0$
Lemma 5.
Lemmas 6 and 7 bound terms appearing in our variance estimators, and will be used to prove Lemma 2.
Lemma 6.
Let Assumption 1 hold, and let $v_{i}$ for $i\in[N]$ denote
If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$, then
If $\phi_{i}^{*}\phi_{j}^{*}=1$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
$\mathbb{E}\left[v_{i},v_{j}|T_{i}T_{j}\right]$ $\leq\frac{1}{p_{\min}^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right].$ (61)
Lemma 7.
Let Assumption 1 hold. Then
$\max_{i:\phi_{i}^{*}=1}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right]$ $=1+O\left(\frac{\log N}{N}\right)$ (62)
$\max_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=1+O\left(\frac{\log N}{N}\right)$ (63)
$\max_{i,j:\phi_{i}^{*}\phi_{j}^{*}=1}{\operatorname{Var}\,}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=O\left(\frac{1}{N}\right)$ (64)
Lemma 8 is used to prove Theorem 6. It gives a coverage result for $\hat{\Phi}_{k}$.
Lemma 8.
Let Assumption 1 hold, and let $\widetilde{V}_{k}$ be defined as in Theorem 6. It holds for $k=1,2$ that
$P\left(\Phi-\hat{\Phi}_{k}\geq z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}\right)\leq\frac{\alpha}{2}+o(1)$
Lemmas 9 and 10 are used to prove Lemma 8. Lemma 9 gives a central limit theorem for a linearized version of $\Phi_{k}$. Lemma 10 is a corollary of Fatou’s lemma, and will be used to show that $V_{k}$ is conservative for the variance of the linearized version of $\Phi_{k}$. It appears as Exercise 3.2.4 in (Durrett, , 2019, 5th edition).
Lemma 9.
Let Assumption 1 hold. For $k\in[2]$, let $\mu_{1}$ and $\mu_{0}$ denote
$\mu_{1}$ $=\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[\tilde{f}_{ki}(T_{i})],$ $\mu_{0}$ $=\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})].$
where $\tilde{f}_{ki}(T_{i})$ for $i:\phi_{i}^{*}=1$ denotes
@@
Let $h_{i}$ for $i\in[N]$ denote
Let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$, and suppose that $\sigma^{2}\geq N^{2/3+\delta}$ for any $\delta>0$. Then $\sum_{i}h_{i}/\sigma$ converges to a standard normal.
Lemma 10.
Theorem 7.
(Ross, , 2011, Thm 3.6) Let $U_{1},\ldots,U_{n}$ be random variables such that $\mathbb{E}[U_{i}^{4}]\leq\infty$, $\mathbb{E}U_{i}=0$, with variance denoted by $\sigma^{2}={\operatorname{Var}\,}\sum_{i=1}^{N}U_{i}$, with dependency neighborhoods $N_{1},\ldots,N_{n}$, which are subsets of $[n]$ (which need not be disjoint), where $U_{j}$ and $U_{i}$ are independent if $U_{j}\notin N_{i}$. Let $D=\max_{i}|N_{i}|$. Let $W=\sum_{i=1}^{N}U_{i}/\sigma$. Then for $Z$ denoting a $N(0,1)$ random variable, and $d_{W}$ denoting Wasserstein metric, it holds that
$d_{W}(W,Z)\leq\frac{D^{3}}{\sigma^{3}}\sum_{i=1}^{n}\mathbb{E}\left[|U_{i}|^{3}\right]+\frac{\sqrt{28}D^{3/2}}{\sqrt{\pi}\cdot\sigma^{2}}\sqrt{\sum_{i=1}^{n}\mathbb{E}\left[X_{i}^{4}\right]}$
A.2 Proof of Theorems
A.2.1 Proof of Theorems 1, 2, and 3
Proof of Theorem 1.
If $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ are within $o_{P}(N)$ of $\Phi$, then $\min\left(\hat{\Phi}_{1}\,,\,\hat{\Phi}_{2}\right)$ is within $o_{P}(N)$ as well. As $\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}(\phi)$ upper bounds $\hat{\Phi}_{k}$, it follows that
$\min\left(\max_{\phi\in\{0,1\}}\hat{\Phi}_{1}(\phi)\,,\,\max_{\phi\in\{0,1\}}\hat{\Phi}_{2}(\phi)\right)\geq\Phi+o_{P}(N),$
which implies that $\hat{\tau}\leq\tau+o_{P}(N)$. ∎
Proof of Theorem 2.
It can be seen that
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}(\phi)$ $=\max_{\phi\in\{0,1\}^{N}}\sum_{i:\phi_{i}=1}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)\phi_{i}$
$=\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)$
$=\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}-\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}+\frac{1-X_{i}}{P(X_{i}=0)}\right)$
$=\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}Y_{i}-\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}\right)+o_{P}(N)$
Where the final equality holds by the assumptions of the theorem. By similar arguments, it holds that
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}(\phi)=\sum_{i=1}^{N}\left(\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}-\frac{X_{i}}{P(X_{i}=1)}Y_{i}\right)+o_{P}(N)$
Substituting into $\hat{\tau}=N-\min\left(\max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{1}(\phi)\ ,\ \max\limits_{\phi\in\{0,1\}}\hat{\Phi}_{2}(\phi)\right)$ yields
$\hat{\tau}=\left|\sum_{i=1}^{N}\frac{X_{i}}{P(X_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}\right|,$
proving the result. ∎
Proof of Theorem 3.
Let $\tilde{f}_{ki}$ be given by (65). Since $\phi_{i}^{*}=1$ and $\bar{Z}_{i}=Y_{ki}$ (which implies $X\in R_{i}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$, and hence that
@@ A.2.1 Proof of Theorems 1, 2, and 3
where (66) holds since $01\{\bar{Z}_{i}=Y_{ki}\}=$ 0 if $X\notin R_{i}$; (67) holds since the $Y_{ki}$ term in the conditional expectation equals $\tilde{f}_{ki}(T_{i})$, and hence may be treated as a constant equal to the observed value of $Y_{ki}$; (68) is a simple rearrangement of terms; and (69) uses that $01\{\bar{Z}_{i}=a\}=$ 0 for $a\in\{0,1\}$ if $X\notin R_{i}$.
By similar reasoning, since $\phi_{i}^{*}\phi_{j}^{*}=1$ and $11\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=$ 1 (which implies that $X\in R_{i}\cap R_{j}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$ and $Y_{kj}=\tilde{f}_{kj}(T_{j})$, and hence that

### v4

@@ C.1 Experiment Designs
    $P(\bar{Z}_{i}=a\,|\,T_{i}=t)$ $\geq p$ $\forall\ i\in[N],a\in\{0,1\},t\in\mathcal{T}_{i}$ (62)
    where $\mathcal{T}_{i}$ is the support of $T_{i}$.
Then Assumption 2 holds. Additionally, Assumption 2 continues to hold if Conditions A3 and A4 are replaced by Condition A6, given below:
  1. A6.
    For $i\in[N]$, $\bar{Z}_{i}$ and $T_{i}$ satisfy
    $\bar{Z}_{i}$ $=\zeta_{i}(\theta_{\eta_{g(i)}})\qquad\textup{ and }$ $T_{i}$ $=X_{G_{g(i)}},$ $\qquad\forall i\in[N],$ (63)
    for some set of mappings $\{\zeta_{i}\}_{i=1}^{N}$.
The conditions Theorem 5 may be interpreted as follows. The conditions (57) and (58) disallows treatment of a vanishing fraction of units, and requires the group level treatment probabilities to be bounded away from zero. The condition (59) requires $\bar{Z}_{i}$ to depend on each unit’s own treatment and the treatment of a bounded number of close associates, and on the number of treated units in their own and neighboring groups. (60) restricts $\mathcal{N}_{i}$ to come from unit $i$’s own and neighboring groups, with size and overlap conditions given by (61). (62) places a support condition on $\bar{Z}_{i}$. (63), which replaces conditions A3 and A4, allows the estimand to count the number of units who are affected by treatments outside of their own group, whose size may increase with $N$.
C.2 Main Theorems
Theorem 6 is a consistency result. It states that $\hat{\tau}^{\textup{Haj}}_{k}$ converges to $\tau$, and that the absolute value of $\Delta^{\textup{Haj}}$, the Hájek weighted contrast between units given by (33), is asymptotically a lower bound for $\tau$. Table 2 gives examples of (33) for the estimands considered in the paper.
Theorem 6.
It holds that
$|\Delta^{\textup{Haj}}|$ $\leq\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ (64)
If Assumption 2 holds, then
$\hat{\tau}_{k}^{\textup{Haj}}=\tau+O_{P}\left(\frac{N}{\sqrt{M}}\right),\qquad k=1,2$ (65)
in which case (64) implies $|\Delta^{\textup{Haj}}|\leq\tau+O_{P}\left(\frac{N}{\sqrt{M}}\right)$
Theorem 7 pertains to variance estimation. It states that $\hat{V}_{k}$ concentrates to its expectation which upper bounds the variance of $\hat{\tau}_{k}^{\textup{Haj}}$, and gives conditions under which the bound will be tight.
Theorem 7.
Let Assumption 2 hold. It holds that
$\hat{V}_{k}=\mathbb{E}[\hat{V}_{k}]+O_{P}\left(\frac{N^{2}\log M}{M^{3/2}}\right),\qquad k=1,2.$ (66)
Additionally, it holds that
$\mathbb{E}\hat{V}_{k}\geq{\operatorname{Var}\,}\hat{\tau}_{k}^{\textup{Haj}},$ (67)
with equality if all of the following conditions hold:
  1. 1.
    The exposure mappings $\{T_{i}\}_{i=1}^{N}$ are constant
  2. 2.
    $P(\bar{Z}_{i}=a)>0$ for all $i\in[N]$ and $a\in\{0,1\}$
  3. 3.
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b)>0$ for all $i,j\in[N]$ such that $i\neq j$, and $a,b\in\{0,1\}$.
Theorem 8 gives coverage properties for the proposed confidence interval given by (36). It states that if the variance estimators $\hat{V}_{k}$ are lower bounded by a vanishing fraction of $\frac{N^{2}}{M}$, then the interval will have at least the desired asymptotic coverage level.
Theorem 8.
@@ C.2 Main Theorems
Let Assumption 2 hold. For any $\epsilon>0$, let $\widetilde{V}_{k}(\phi)$ denote the thresholded variance estimate given by
$\widetilde{V}_{k}(\phi)=\max\left\{\hat{V}_{k}(\phi)\ ,\ \frac{1}{z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2}\cdot\frac{N^{2}}{M^{4/3-\epsilon}}\right\}$ (68)
and let $LB_{1-\alpha}$ denote
$LB_{1-\alpha}=\max\left\{\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{1}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\,,\,\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{2}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}$ (69)
Then $\tau\geq LB_{1-\alpha}$ with probability converging to at least $1-\alpha$.
Remark
Theorems 1 - 3 can be seen to follow directly from Theorems 6 - 8, by substituting $M=N$ and applying Theorem 4.
Remark
The rates appearing in Theorems 6 - 8 may be more familiar or intuitive after normalizing the estimand by $N$, so that it equals $1-|{\mathcal{I}}|/N$, the fraction of units not in $\mathcal{I}$}, and then normalizing $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$ accordingly by $N$ and $N^{2}$ accordingly. For this choice of normalization, Theorem 6 implies $O_{P}(1/\sqrt{M})$ estimation error, as might be expected for an experiment whose randomization was constructed from $M$ independent draws. Similarly, Theorem 7 implies $o_{P}(1/M)$ error for the variance estimator after normalization by $N$, with Theorem 8 lower bounding the estimated variance by $o_{P}(1/M)$ as well.
Appendix D Proofs
Propositions 1 - 5 are proven in Section D.1. Theorems 4 and 5 are proven in Section D.2. Proofs of Theorems 6, 7, and 8 (which are generalized versions of Theorems 1, 2, and 3), including auxilliary lemmas, are respectively contained in Sections D.3, D.4, and D.5.
D.1 Proof of Propositions 1 - 5
Proof of Proposition 1.
It can be seen that
$\hat{\tau}_{1}$ $=N-\sum_{i\in{\mathcal{I}}}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)$ (70)
$\geq N-\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})(1-Y_{i})}{P(X_{i}=0)}\right)$ (71)
$=N-\sum_{i=1}^{N}\left(\frac{X_{i}Y_{i}}{P(X_{i}=1)}-\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}+\frac{1-X_{i}}{P(X_{i}=0)}\right)$ (72)
$=\sum_{i=1}^{N}\left(\frac{1-X_{i}}{P(X_{i}=0)}-\frac{X_{i}}{P(X_{i}=1)}\right)Y_{i}+O_{P}(N^{1/2})$ (73)
where (70) uses the definition of $\hat{\tau}_{1}$ given in (5) and that
$P(H_{i}=1)=Y_{i}P(X_{i}=1)+(1-Y_{i})P(X_{i}=0)\qquad\forall\ i\in{\mathcal{I}}$
as $Y_{i}$ is constant if $i\in{\mathcal{I}}$; (71) holds as the terms in the summand are nonnegative; (72) is simple rearrangement of terms, and (73) uses (6) assumed by the lemma.
By essentially identical arguments, it holds that
$\hat{\tau}_{2}$ $=N-\sum_{i\in{\mathcal{I}}}\left(\frac{X_{i}(1-Y_{i})}{P(X_{i}=1)}+\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}\right)$
$\geq N-\sum_{i=1}^{N}\left(\frac{X_{i}(1-Y_{i})}{P(X_{i}=1)}+\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}\right)$
$\geq N-\sum_{i=1}^{N}\left(-\frac{X_{i}Y_{i}}{P(X_{i}=1)}+\frac{(1-X_{i})Y_{i}}{P(X_{i}=0)}+\frac{X_{i}}{P(X_{i}=1)}\right)$
$=\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}+O_{P}(N^{1/2})$ (74)
Combining (73) and (74) yields
$\max(\hat{\tau}_{1},\hat{\tau}_{2})\geq\left|\sum_{i=1}^{N}\left(\frac{X_{i}}{P(X_{i}=1)}-\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}\right|,$
proving the result. ∎
Proof of Proposition 2.
Substitute $Z_{i}$ for $X_{i}$ in the proof of Theorem 1. ∎
Proof of Proposition 3.
By (28) and the conditions of the proposition, it holds for $a\in\{0,1\}$ that
$1\{\bar{Z}_{i}=a\}Y_{i}=\begin{cases}1\{\bar{Z}_{i}=a\}\tilde{f}_{i}(T_{i})&\textup{ if }i\in{\mathcal{I}}\\ 1\{\bar{Z}_{i}=a\}\bar{Z}_{i}&\textup{ if }i\notin{\mathcal{I}}\end{cases}$ (75)
where we have used that $\bar{Z}_{i}\in\{0,1\}$ implies $X\in R_{i}$ by (29). Plugging (75) into $\Delta^{\bar{Z}}$ yields
$\Delta^{\bar{Z}}$ $=\sum_{i\in{\mathcal{I}}}\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)\tilde{f}_{i}(T_{i})+\sum_{i\notin{\mathcal{I}}}\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)\bar{Z}_{i}$
It can be seen that the first summation (which is over $i\in{\mathcal{I}}$) has expectation zero, while the remaining summation has expectation $|\{i:i\notin{\mathcal{I}}\}|=N-|{\mathcal{I}}|$, proving the proposition. ∎
Proof of Proposition 4.
Let $\tilde{f}_{ki}$ for $i\in{\mathcal{I}}$ denote the mapping
@@ D.1 Proof of Propositions 1 - 5
As $Y_{ki}=\tilde{f}_{ki}(T_{i})$ whenever $X\in R_{i}$ and $i\in{\mathcal{I}}$, it follows that
$P(X\in R_{i},Z_{i}=Y_{ki}|T_{i})$ $=P(X\in R_{i},Z_{i}=\tilde{f}_{ki}(T_{i})|T_{i})$ (77)
$=\sum_{a=0}^{1}1\{\tilde{f}_{ki}(T_{i})=a\}\ P(X\in R_{i},Z_{i}=a|T_{i},X\in R_{i})$
$=\tilde{f}_{ki}(T_{i})\cdot P(X\in R_{i},Z_{i}=1|T_{i})+(1-\tilde{f}_{ki}(T_{i}))\cdot P(X\in R_{i},Z_{i}=0|T_{i})$
$=Y_{ki}\cdot P(X\in R_{i},Z_{i}=1|T_{i})+(1-Y_{ki})\cdot P(X\in R_{i},Z_{i}=0|T_{i})$ (78)
where (77) holds because $i\in{\mathcal{I}}$, and (78) holds because $X\in R_{i}$ and $i\in{\mathcal{I}}$, so that we $Y_{ki}=\tilde{f}_{ki}(T_{i})$
Proof of Proposition 5.
Let $\tilde{f}_{ki}$ be given by (76). Since $i\in{\mathcal{I}}$ and $\bar{Z}_{i}=Y_{ki}$ (which implies $X\in R_{i}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$, and hence that
@@
where (79) holds since $1\{\bar{Z}_{i}=Y_{ki}\}=0$ if $X\notin R_{i}$; (80) holds since the $Y_{ki}$ term in the conditional expectation equals $\tilde{f}_{ki}(T_{i})$, and hence may be treated as a constant equal to the observed value of $Y_{ki}$; (81) is a simple rearrangement of terms; and (82) uses that $1\{\bar{Z}_{i}=a\}=0$ for $a\in\{0,1\}$ if $X\notin R_{i}$.
By similar reasoning, if $i,j\in{\mathcal{I}}$ and $1\{\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\}=1$ (which implies that $X\in R_{i}\cap R_{j}$), it holds that $Y_{ki}=\tilde{f}_{ki}(T_{i})$ and $Y_{kj}=\tilde{f}_{kj}(T_{j})$, and hence that